OpenStudy (kordel):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent? (4 points)

OpenStudy (kordel):

someon please help

OpenStudy (mathstudent55):

Do you know what the domain of a function is?

OpenStudy (wckeller):

Part A: To find a reasonable domain, set the y value to the maximum height of the plant. 11.26=8(1.05)n n=11.26/(8(1.05)) n=1.34 Because the plant cannot be a negative height, our minimum for the domain must be 0. The domain of the function should be [0,1.34] Part B: The Y intercept of the function represents the time, in days, when the plant had a height of 0. Part C: Average rate of change, A.K.A the slope of the function. is 8*1.05=8.4 This represents how much the plant grew each day. (8.4cm)

OpenStudy (mathstudent55):

The function is written above as f(n) = 8(1.05)n. That makes no sense. If the 8 and the 1.05 and the n are just factors in a multiplication, then it could easily be written as f(n) = 8.4n Is the function really \(f(n) = 8(1.05)^n\) ? If so, get used to writing exponents properly because O.S. helpers can't guess what you mean, and then just waste time trying to help you with an incorrect equation.